OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

i plugged them in but i got a weird answer

OpenStudy (anonymous):

-5=2(1)+4

OpenStudy (anonymous):

Since they r parallel Slopes are =

OpenStudy (anonymous):

ie m1 = m2

OpenStudy (anonymous):

coption is right

OpenStudy (anonymous):

y = 2x + 4. Slope = 2

OpenStudy (anonymous):

(1, -5) y + 5 = 2(x-1) y + 5 = 2x - 2 y = 2x - 7

OpenStudy (ajprincess):

slope=2 since slopes are same for parallel lines. \(y-y_1=m(x-x_1)\) \(y-(-5)=2(x-1)\) solve for y and u will get the answer. Can u do it @carachronic?